# Pro-Pose: Unpaired Full-Body Portrait Synthesis via Canonical UV Maps

### v1

@@ 8 Canonical UV-Space Formulation
Figure 11: Random Patch Masking Augmentation vs. Ours (Donor-Based UV Reposing). We compare random patch masking texture augmentation against our donor-based UV reposing method. With standard random masking, the shape of the UV boundaries inadvertently reveals pose information. This leakage causes the model to generate significant artifacts, particularly near face boundaries. As seen in the first row, hand-face occlusions create hand artifacts, while the second and third rows display hallucinated accessories such as swimming goggles, hats, and dreadlocks. By effectively decoupling texture from pose via donor-based UV reposing, our method resolves these issues, yielding cleaner results with superior identity preservation.
@@
9 Ablation Study: Donor-based UV Reposing vs Random Patch Masking
In Section 3.2.1 (main paper), we introduced Donor-based UV Reposing to prevent the model from learning shortcuts. We argued that using a real mask from a ”donor” creates realistic occlusions that force the network to learn geometry. In this section, we test if this is actually necessary. We compare our method against a simpler baseline where we just block out random parts of the texture to create a synthetic pair.
Implementation Details. For the ”Random masking” baseline, we take the input facial texture ($\mathbf{T}_{\mathbf{p}}$) and cover it with random black squares. We use up to 6 non-overlapping patches, each $64\times 64$ pixels. All other training settings remain identical to our main method.
Qualitative Results. Visual comparisons in Figure 11 show that random masking is insufficient. The problem is that random squares do not change the overall ”shape” of the visible texture. The boundaries of the valid texture still look like the source pose. Because the model can still recognize the source pose from these boundaries, it tries to find a shortcut. However, the random black squares create confusing holes that don’t match natural occlusions. To explain these unnatural missing regions, the model, more frequently, hallucinates accessories.
Our Donor-based method avoids this because the donor mask looks like a real human pose. It changes the texture boundaries completely, so the model cannot rely on the visible pose outline and must learn beyond simple ”copy-pasting”.
@@ 9 Ablation Study: Donor-based UV Reposing vs Random Patch Masking
Random masking 18.56 6.45 0.821 0.5187 0.6965 7.35
Ours 19.87 6.04 0.849 0.5563 0.7383 7.49
Table 3: Donor Masking Ablation. We compare our method against a ”Random Masking” baseline, that applies multiple black patches ($64\times 64$) to the input texture. The superior performance of our method confirms that structurally semantic donor masks are essential for learning robust geometry, whereas random patches fail to prevent boundary-based pose leakage.
10 Additional Results
@@
Figure 12: Qualitative Comparison. The last three rows display results from the DeepFashion dataset [25], and all previous rows utilize real images from the WPose dataset [20]. The columns show the Original Person, Target Pose, and results from state-of-the-art methods (MCLD [24], Leffa [53], UniHuman [20], OnePoseTrans [6]) compared to Ours and the Ground Truth (GT). Our method demonstrates superior performance in preserving the person’s identity, particularly the facial characteristics, across varying poses. For instance, in the third and fifth rows, previous methods introduce significant identity distortion and loss of facial likeness (e.g., changes in jawline or features). In contrast, Ours robustly preserves the unique face structure and identity of the original person, closely matching the GT result.
@@
CFLD [26] 14.34 0.7012 0.1823 0.9543
MCLD [24] 15.76 0.7151 0.1798 0.9429
LEFFA [53] 13.95 0.6956 0.12 0.9316
OnePoseTrans [6] 13.12 0.5738 0.322 0.9319
UniHuman [20] 13.87 0.7557 0.161 0.9534
@@
CFLD [26] 15.12 0.741 0.279 0.6109
MCLD [24] 15.29 0.757 0.24 0.6108
LEFFA [53] 16.47 0.77 0.236 0.5541
OnePoseTrans [6] 17.1 0.807 0.174 0.6719
UniHuman [20] 17.49 0.802 0.179 0.6814
@@
Table 4: Quantitative Evaluation and Ablation Study. We compute PSNR, SSIM, LPIPS and DINO only on the face region of the generated images.
10.1 Extended Evaluation In-The-Wild
To demonstrate the robustness of our approach, we evaluate Pro-Pose on “in-the-wild” images sourced from CelebA-HQ [15]. These samples feature diverse lighting conditions, backgrounds, and subjects. The qualitative results, presented in Figure 10, confirm that our method generalizes effectively to unseen real-world data, maintaining high identity fidelity and pose accuracy.
10.2 Extended Qualitative Comparisons
We provide more extensive qualitative comparisons against state-of-the-art methods (MCLD [24], Leffa [53], UniHuman [20] , OnePoseTrans [6]) in Figure 12. Our method consistently produces higher fidelity faces and fewer geometric artifacts than competing approaches, particularly for challenging poses and diverse identities. Our method performs equally well on both in-domain (DeepFashion [25]) and in-the-wild (WPose [20]) scenarios, maintaining consistent quality across datasets.
10.3 Extended Quantitative Evaluation
To further validate the fidelity of our generated avatars, we report metrics computed exclusively on the facial regions in Table 4. This localized evaluation prevents the large, uniform areas of background and body from inflating the pixel-level metrics, thereby providing a stricter assessment of identity and facial detail preservation.
@@ 10.3 Extended Quantitative Evaluation
Analysis. The results reinforce the generalization gap observed in our main experiments. On the in-domain DeepFashion dataset, the ”Paired Only” model performs marginally better, consistent with its tendency to overfit to the training distribution. However, on the challenging, out-of-distribution WPose dataset, Pro-Pose significantly outperforms all baselines and the ”Paired Only” variant across all metrics. This confirms that our hybrid training strategy yields superior generalization on fine-grained facial details, avoiding the identity degradation observed in baselines.
10.4 Impact of Training Data Source
Complementing the quantitative ablation in the main paper, Figure 13 visually demonstrates the impact of our hybrid data strategy. Models trained on ”Unpaired Only” data struggle with extreme poses, while ”Paired Only” models suffer from identity drift. Our combined approach (”Unpaired + Paired”) leverages the strengths of both, achieving robust pose control and identity preservation.
10.5 Personalization with Few-Shot Learning
@@
More Qualitative Evaluation. We demonstrate how our few-shot learning strategy generalizes to in-the-wild scenarios [54] in Figure 14. Notably, the identities of the generated new poses become significantly closer to the original subject after fine-tuning, validating the effectiveness of this adaptation step.
Figure 15: DeepFashion. We highlight the top 100 identity clusters found in the DeepFashion dataset. We also exclusively highlight the first person on the top-left, who can be repeatedly obersved in the generations of the overfitted models.
@@
11 Analysis of Identity Overfitting
@@
11.1 DeepFashion Identity Analysis
To empirically validate the limited identity diversity of the DeepFashion In-Shop Clothes Retrieval benchmark [25], we performed a graph-based clustering analysis on the training set. Our pipeline proceeds as follows:
@@ 11.1 DeepFashion Identity Analysis
    Embedding Extraction: We utilized ArcFace [4] to detect faces and extract identity embeddings. To ensure embedding reliability, we explicitly filtered out faces with extreme head poses (yaw or pitch $45^{\circ}$).
@@
11.2 Identity Analysis of Generated Images
We diagnose this phenomenon in Figure 16. We reposed subjects from the out-of-distribution - WPose dataset using two models: one trained only on paired DeepFashion data, and our full model trained on both paired and unpaired data. We then measured the face similarity between the generated outputs and the DeepFashion training set.
  • The Paired-Only model (red histograms in Figure 16 (a) and (b)) produces outputs with high similarity to the training data, indicating it is ”pulling” novel identities toward the few identities it memorized during training.
@@ 11.2 Identity Analysis of Generated Images
11.3 DeepFashion Identity Limitation Leading to Overfitting
We highlight a dominant identity from the DeepFashion training set in the red box of Figure 15. Comparing this to the ”Paired Only” results in Figure 13 (and Figure 6 of main paper) reveals severe overfitting: As we can see in some cases the model tends to collapse to this specific training identity (for example second row in Figure 13). This bias is so strong that it overrides semantic gender cues, frequently substituting female subjects with this specific male face, confirming that the paired-only model relies on memorization rather than generalization.

### v2

@@ Appendix 0.B Canonical UV-Space Formulation
Appendix 0.C Ablation Study: Donor-based UV Reposing vs Random Patch Masking
Figure 13: Random Patch Masking vs. Ours (Donor-Based UV Reposing). Random masking preserves UV boundary shape, leaking pose information and causing artifacts (e.g., hallucinated accessories). Our donor-based reposing changes texture boundaries completely, forcing the model to learn geometry rather than shortcuts.
In Section 3.2 (main paper), we introduced Donor-based UV Reposing to prevent the model from learning shortcuts. We argued that using a real mask from a "donor" creates realistic occlusions that force the network to learn geometry. In this section, we test if this is actually necessary. We compare our method against a simpler baseline where we just block out random parts of the texture to create a synthetic pair.
Implementation Details. For the "Random masking" baseline, we take the input texture ($\mathbf{T}_{\mathbf{p}}$) and cover it with random black squares. We use up to 6 non-overlapping patches, each $64\times 64$ pixels. All other training settings remain identical to our main method.
Qualitative Results. Visual comparisons in Figure 13 show that random masking is insufficient. The problem is that random squares do not change the overall "shape" of the visible texture. The boundaries of the valid texture still look like the source pose. Because the model can still recognize the source pose from these boundaries, it tries to find a shortcut. However, the random black squares create confusing holes that do not match natural occlusions. To explain these unnatural missing regions, the model, more frequently, hallucinates accessories.
Our Donor-based method avoids this because the donor mask looks like a real human pose. It changes the texture boundaries completely, so the model cannot rely on the visible pose outline and must learn beyond simple "copy-pasting".
[…]
Random 18.56 6.45 0.821 0.5187 0.6965 7.35
Ours 19.87 6.04 0.849 0.5563 0.7383 7.49
Table 3: Donor Masking Ablation. Random masking ($64\times 64$ patches) vs. our donor-based reposing. Our method outperforms on all metrics, confirming that structurally semantic masks are essential.
@@ Appendix 0.C Ablation Study: Donor-based UV Reposing vs Random Patch Masking
Appendix 0.D Additional Results
DeepFashion (In-Domain) – BC Inputs
@@ Appendix 0.D Additional Results
CFLD [CFLD] 18.51 7.09 0.757 0.180 0.48 0.3182 0.9733 4.21
MCLD [MCLD] 19.18 7.01 0.769 0.173 0.49 0.3437 0.9657 4.35
LEFFA [Zhou2025LEFFA] 14.95 4.20 0.767 0.117 0.45 0.5781 0.9415 4.40
OnePoseTrans [oneposetrans] 14.35 8.59 0.618 0.287 0.46 0.5761 0.9481 4.37
UniHuman [Li_UniHuman_2024] 15.95 6.16 0.803 0.149 0.46 0.5805 0.9436 4.24
Gemini 2.5 Flash Image [nanobanana] 17.59 4.44 0.757 0.167 0.43 0.5817 0.9697 7.22
Ours 19.36 4.24 0.818 0.075 0.48 0.6047 0.9759 7.24
WPose (Out-of-Domain) – BC Inputs
[…]
CFLD [CFLD] 15.61 91.67 0.753 0.201 0.31 0.0888 0.6418 1.92
MCLD [MCLD] 15.80 88.76 0.764 0.196 0.35 0.0997 0.6479 1.93
LEFFA [Zhou2025LEFFA] 16.82 64.44 0.786 0.186 0.33 0.0911 0.5734 2.04
OnePoseTrans [oneposetrans] 17.35 25.87 0.825 0.145 0.33 0.1734 0.7214 4.39
UniHuman [Li_UniHuman_2024] 17.75 24.13 0.815 0.153 0.34 0.1121 0.7211 2.95
Gemini 2.5 Flash Image [nanobanana] 16.74 8.18 0.788 0.144 0.32 0.4727 0.7021 7.38
Ours 19.95 5.99 0.860 0.121 0.38 0.5571 0.7394 7.55
Table 4: Quantitative Comparison with BC Inputs. All baselines receive BC-preprocessed input images at test time, removing garment variation as a confounding factor. Pro-Pose maintains its advantage, confirming that our gains are methodological.
Figure 14: Pro-Pose In-the-wild. We apply our method to real-world images characterized by varying poses and environments. The columns display (from left to right): (1) the original input image; (2) the target pose represented by a SMPL-X mesh (reposed using ground truth parameters); (3) the output generated by Pro-Pose with a corresponding face crop for detailed visual comparison. Our method consistently preserves identity, including facial details, body shape, and visible skin features, compared to the ground truth.
0.D.1 Extended Evaluation In-The-Wild
To demonstrate the robustness of our approach, we evaluate Pro-Pose on “in-the-wild” images sourced from CelebA-HQ [celebAHQ]. These samples feature diverse lighting conditions, backgrounds, and subjects. The qualitative results, presented in Figure 14, confirm that our method generalizes effectively to unseen real-world data, maintaining high identity fidelity and pose accuracy.
@@ 0.D.1 Extended Evaluation In-The-Wild
Figure 15: Qualitative Comparison. The last three rows display results from the DeepFashion dataset [deepfashion], and all previous rows utilize real images from the WPose dataset [Li_UniHuman_2024]. The columns show the Original Person, Target Pose, and results from state-of-the-art methods (MCLD [MCLD], Leffa [Zhou2025LEFFA], UniHuman [Li_UniHuman_2024], OnePoseTrans [oneposetrans], Gemini 2.5 Flash [nanobanana]) compared to Ours and the Ground Truth (GT). Our method demonstrates superior performance in preserving the person’s identity, including facial features, body shape, and visible skin characteristics, across varying poses. For instance, in the third and fifth rows, previous methods introduce significant identity distortion and loss of facial likeness (e.g., changes in jawline or features). In contrast, Ours robustly preserves the unique face structure and identity of the original person, closely matching the GT result. For the final row, no ground truth (GT) exists because the input is a single image reposed into a target A-pose. Although a GT face crop cannot be provided, comparing the result to the face crop of the original image demonstrates our strong identity preservation.
0.D.2 Extended Qualitative Comparisons
We provide more extensive qualitative comparisons against state-of-the-art methods (MCLD [MCLD], Leffa [Zhou2025LEFFA], UniHuman [Li_UniHuman_2024], OnePoseTrans [oneposetrans], and Gemini 2.5 Flash [nanobanana]) in Figure 15. Our method consistently produces higher fidelity identity preservation and fewer geometric artifacts than competing approaches, particularly for challenging poses and diverse identities. Our method performs equally well on both in-domain (DeepFashion [deepfashion]) and in-the-wild (WPose [Li_UniHuman_2024]) scenarios, maintaining consistent quality across datasets.
0.D.3 Evaluation with BC Inputs for Baselines
A key difference between Pro-Pose and prior methods is that our model is trained on Base Clothing (BC) standardized inputs, while all baselines are trained on original images with diverse garments. To ensure a fair comparison, we additionally evaluate all baselines when provided with BC-preprocessed input images at test time. This removes garment variation as a confounding factor and isolates each method’s ability to preserve identity and follow the target pose.
Quantitative Results. Table 4 reports the results. Even when baselines receive the same BC inputs that align with our training distribution, Pro-Pose maintains a significant advantage across identity (FaceSim) and perceptual quality (HPSv3) metrics, confirming that our gains stem from the method itself rather than the data preprocessing.
Qualitative Results. Figure 16 shows representative examples. Despite receiving the same standardized inputs, baselines still exhibit characteristic failures: identity drift (CFLD, MCLD), geometric artifacts in extreme poses (OnePoseTrans), and over-smoothed facial details (UniHuman). Pro-Pose consistently produces sharper, more identity-faithful results.
0.D.4 Extended Quantitative Evaluation
@@
CFLD [CFLD] 14.34 0.7012 0.1823 0.9543
MCLD [MCLD] 15.76 0.7151 0.1798 0.9429
LEFFA [Zhou2025LEFFA] 13.95 0.6956 0.12 0.9316
OnePoseTrans [oneposetrans] 13.12 0.5738 0.322 0.9319
UniHuman [Li_UniHuman_2024] 13.87 0.7557 0.161 0.9534
Gemini 2.5 Flash [nanobanana] 15.59 0.7107 0.1018 0.9531
@@
CFLD [CFLD] 15.12 0.741 0.279 0.6109
MCLD [MCLD] 15.29 0.757 0.24 0.6108
LEFFA [Zhou2025LEFFA] 16.47 0.77 0.236 0.5541
OnePoseTrans [oneposetrans] 17.1 0.807 0.174 0.6719
UniHuman [Li_UniHuman_2024] 17.49 0.802 0.179 0.6814
Gemini 2.5 Flash [nanobanana] 16.96 0.781 0.198 0.6776
@@
Table 5: Extended Quantitative Evaluation. Complementing the full-body metrics in the main paper, we compute PSNR, SSIM, LPIPS, and DINO for the face region.
To further validate the fidelity of our generated avatars, we report metrics computed exclusively on the facial regions in Table 5. This complements the full-body metrics in the main paper by providing a stricter assessment of identity and facial detail preservation, where the large, uniform areas of background and body cannot inflate pixel-level scores.
@@
Analysis. The results reinforce the generalization gap observed in our main experiments. On the in-domain DeepFashion dataset, the "Paired Only" model performs marginally better, consistent with its tendency to overfit to the training distribution. However, on the challenging, out-of-distribution WPose dataset, Pro-Pose significantly outperforms all baselines and the "Paired Only" variant across all metrics. This confirms that our hybrid training strategy yields superior generalization on fine-grained identity features, avoiding the identity degradation observed in baselines.
Figure 16: Qualitative Comparison with BC Inputs. All methods receive BC-preprocessed images as input, isolating the effect of the generation method from the data preprocessing. Pro-Pose preserves identity and pose fidelity more consistently than all baselines. For the final row, no ground truth (GT) exists because the input is a single image reposed into a target A-pose. Although a GT face crop cannot be provided, comparing the result to the face crop of the original image demonstrates our strong identity preservation.
0.D.5 Impact of Training Data Source
Complementing the quantitative ablation in the main paper, Figure 17 visually demonstrates the impact of our hybrid data strategy. Models trained on "Unpaired Only" data struggle with extreme poses, while "Paired Only" models suffer from identity drift. Our combined approach ("Unpaired + Paired") leverages the strengths of both, achieving robust pose control and identity preservation.
0.D.6 Personalization with Few-Shot Learning
@@
More Qualitative Evaluation. We demonstrate how our few-shot learning strategy generalizes to in-the-wild scenarios [celebVHQ] in Figure 18. Notably, the identities of the generated new poses become significantly closer to the original subject after fine-tuning, validating the effectiveness of this adaptation step.
Figure 19: DeepFashion. We highlight the top 100 identity clusters found in the DeepFashion dataset. We also exclusively highlight the first person on the top-left, who can be repeatedly observed in the generations of the overfitted models.
@@
Appendix 0.E Analysis of Identity Overfitting
@@
0.E.1 DeepFashion Identity Analysis
To empirically validate the limited identity diversity of the DeepFashion In-Shop Clothes Retrieval benchmark [deepfashion], we performed a graph-based clustering analysis on the training set. Our pipeline proceeds as follows:
@@ 0.E.1 DeepFashion Identity Analysis
    Embedding Extraction: We utilized ArcFace [arcface] to detect faces and extract identity embeddings. To ensure embedding reliability, we explicitly filtered out faces with extreme head poses (yaw or pitch $45^{\circ}$).
@@
0.E.2 Identity Analysis of Generated Images
We diagnose this phenomenon in Figure 20. We reposed subjects from the out-of-distribution WPose dataset using two models: one trained only on paired DeepFashion data, and our full model trained on both paired and unpaired data. We then measured the face similarity between the generated outputs and the DeepFashion training set.
  • The Paired-Only model (red histograms in Figure 20 (a) and (b)) produces outputs with high similarity to the training data, indicating it is "pulling" novel identities toward the few identities it memorized during training.
@@ 0.E.2 Identity Analysis of Generated Images
0.E.3 DeepFashion Identity Limitation Leading to Overfitting
We highlight a dominant identity from the DeepFashion training set in the red box of Figure 19. Comparing this to the "Paired Only" results in Figure 17 (and Figure 7 of main paper) reveals severe overfitting: as we can see, in some cases the model tends to collapse to this specific training identity (for example second row in Figure 17). This bias is so strong that it overrides semantic gender cues, frequently substituting female subjects with this specific male face, confirming that the paired-only model relies on memorization rather than generalization.